OpenStudy (anonymous):

To graph the function g(x)=-(x-4)^3, start with the graph f(x)=x^3 then make which of the following changes? A. Shift the graph of left by 4 units. B. Shift the graph of right by 4 units. C. Shift the graph of up by 4 units. D. Shift the graph of down by 4 units. E. Reflect the graph of over the x-axis. F. Reflect the graph of over the y-axis. G. Stretch the graph of horizontally. H. Stretch the graph of vertically.

OpenStudy (nirmalnema):

Shift the graph of right by 4 units.

OpenStudy (anonymous):

Is that all you do?

OpenStudy (nirmalnema):

Reflect the graph of over the y-axis. due to presence of - sign

OpenStudy (anonymous):

So B and F? That appears to be wrong

OpenStudy (phi):

if you have y= 2 (for example) and then multiply the right side by -1 you get y=-2 you "flipped" over the X-axis.

OpenStudy (anonymous):

Then what would it be for something like x^3+7 in relation to g(x)=X^3

OpenStudy (phi):

x^3 + 7 is the same as x^3 (except obviously we add +7) if you add +7 you make the graph shift up

OpenStudy (anonymous):

Okay. Thank you very much!